Two identical spheres each of mass 1.2 kg and radius 10 centimetre are fixed at the end of a light rod so that the separation between the centre is 50 cm find the moment of inertia about an Axis perpendicular to the rod passing through its middle.

Grade:12

1 Answers

jitender
6 years ago
I = (mr^2)/12 Given m = 1.2 each so total M=2.4kgR=50cm =0.5mI = (2.4×0.5^2)/12=2.4×0.25/12=0.2×0.25 =0.050 Kg m2